# PoliPrompt: A High-Performance Cost-Effective LLM-Based Text Classification Framework for Political Science

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have opened new avenues for enhancing text classification efficiency in political science, surpassing traditional machine learning methods that often require extensive feature engineering, human labeling, and task-specific training. However, their effectiveness in achieving high classification accuracy remains questionable. This paper introduces a three-stage in-context learning approach that leverages LLMs to improve classification accuracy while minimizing experimental costs. Our method incorporates automatic enhanced prompt generation, adaptive exemplar selection, and a consensus mechanism that resolves discrepancies between two weaker LLMs, refined by an advanced LLM. We validate our approach using datasets from the BBC news reports, Kavanaugh Supreme Court confirmation, and 2018 election campaign ads. The results show significant improvements in classification F1 score (+0.36 for zero-shot classification) with manageable economic costs (-78% compared with human labeling), demonstrating that our method effectively addresses the limitations of traditional machine learning while offering a scalable and reliable solution for text analysis in political science.
@@
text classification, sentiment analysis, large language models (LLMs)
@@ 1 Introduction
Text has always been an important data source in political science. Text analysis in political science can be categorized into several key tasks, including classification (boussalis2016text; farrell2016corporate), scaling (barbera2015understanding; lauderdale2016measuring), text reuse detection (e.g., hertel2015capturing; hertel2018politics), and natural language processing (e.g., bird2009natural; leetaru2013gdelt). Among these, classification is the most common task in text analysis in the field. For instance, researchers might investigate whether campaign ads are positive or negative across different media platforms (fowler2021political), or examine if online posts and newspaper coverage in authoritarian countries highlight local government wrongdoing (Pan and Chen 2018). These examples, along with many others in political science, demonstrate that understanding the nature of politics requires insight into what political actors are saying and writing. With recent advances in theory and practice, political scientists increasingly rely on machine learning (ML) methods to classify large corpora of text by measures such as topic and tone (grimmer2013text).
Figure 1: The overview of the three-stage framework with LLM.
Traditional machine learning methods (8950616; 9079815; 1174243; devlin2019bertpretrainingdeepbidirectional; sun2020finetuneberttextclassification), such as Bag of Words (BoW), Support Vector Machines (SVM), and Bidirectional Encoder Representations from Transformers (BERT), have significantly contributed to text classification. Models like BoW and SVM rely on feature extraction from text or selection of features from correlated data, often requiring manual tuning and domain expertise for optimal performance. While BERT represents a more advanced approach, it demands substantial computational resources and extensive data for fine-tuning. These methods necessitate task-specific training, meaning that each new task requires a fresh round of model training to achieve optimal results. For example, transitioning from topic modeling to sentiment analysis requires training an entirely new model. As a result, these approaches typically rely heavily on human annotators to label data for training, making them labor-intensive, costly, and less flexible. This lack of generalizability across tasks is a major limitation, reducing the reusability of the model for different text classification challenges and piling up the cost.
In recent years, large language models (LLMs) such as Mistral (jiang2024mixtralexperts) and GPT (Radford2018ImprovingLU) have emerged as powerful tools for text analysis, offering the ability to generate predictions without the need for task-specific training. These models, pre-trained on a vast corpus of text, possess a deep understanding of language and can perform zero-shot and few-shot predictions—wherein the model makes predictions without task-specific training [^1] For instance, unlike traditional machine learning methods that require task-specific training sets—where humans must label data according to the task at hand, such as support/oppose for stance or negative/neutral/positive for sentiment—LLM eliminates the need for such labor-intensive model training each time the task changes.. This capability represents a significant leap forward in terms of efficiency and accessibility, particularly for social scientists who may lack the resources to develop and train their own models.
However, despite their advantages, LLMs are not without their own set of challenges. One of the primary drawbacks is that they can produce inaccurate predictions when provided with simple or poorly constructed prompts (brown2020languagemodelsfewshotlearners). Even when incorporating a few-shot learning approach—wherein the model is given a very small number of annotated examples to guide its predictions—the fixed nature of these examples may result in poor performance if the examples do not adequately represent the broader dataset. Besides, current practices (Wang_2023; egami2024using) that only provide a label obscure the reasoning behind the LLM’s predictions from humans. As a result, if the LLM misinterprets the human-constructed prompt, this misunderstanding goes unnoticed, limiting the opportunity for human-LLM interaction to refine the prompt, inject additional human knowledge, and perform validation. Moreover, while advanced LLMs such as GPT-4 offer higher levels of accuracy and general intelligence, they come with significant costs and usage restrictions compared with weak ones such as GPT-3.5 (Table 1).
Feature GPT-4-turbo GPT-3.5-turbo
Input Token Pricing \$10 / 1M tokens \$0.5 / 1M tokens
Output Token Pricing \$30 / 1M tokens \$1.5 / 1M tokens
Tokens per Minute 10,000 200,000
Batch Queue Limit 100,000 2,000,000
Table 1: Comparison of GPT-4-turbo and GPT-3.5-turbo at the time of writing (openai_pricing).
To address these challenges, we propose a novel LLM-based three-stage (pre-process, in-process, post-process) framework (Figure 1) for text classification which achieves high accuracy without the need for task-specific training and requires only minimal human annotations.
First of all, our approach selects a small pool of diverse examples to represent the broader dataset, guiding the LLM to extract classification rules and write enhanced prompts based on accurate human-annotated texts. During inference, it dynamically selects examples that are most similar to the query text [^2] A query text is the text a user asks the LLM to annotate., thereby enhancing the relevance of the examples used for prediction. This dynamic selection process addresses the limitations of fixed examples in few-shot learning and significantly improves the accuracy of the model’s predictions.
Furthermore, our method capitalizes on the consensus among two weaker LLMs, rather than relying on a single advanced model. By aggregating the outputs of these models, we achieve a level of accuracy comparable to that of more sophisticated LLMs at a fraction of the cost. This approach not only reduces computational time but also circumvents the restrictive usage limits imposed by the service provider of the advanced model.
Another key advantage of our framework is its ability to provide explanations for the model’s predictions. Unlike traditional machine learning methods that only give a prediction, our method produces interpretable reasoning behind each annotation for hard queries. This transparency is crucial in social sciences, where the stakes of misinterpretation are high, and the ability to understand and justify the reasoning behind predictions is essential for scholarly rigor.
In the following sections, we will delve into our methodology and experiments, elaborate on the implementation details, clarify the rationale behind each module, validate the outlined advantages, and demonstrate the framework’s effectiveness through text classification tasks in political science.
@@ 2 Literature Review
Supervised machine learning for text classification.
Traditional machine learning methods such as BoW, N-grams, and TF-IDF combined with SVM or Bayes have played a significant role in the analysis of text data within the social sciences (collingwood2012tradeoffs; drutman2013inside; grimmer2013text; ceron2014every; wilkerson2017large). These methods, however, come with substantial limitations. At their core, traditional supervised machine learning approaches require intensive feature engineering—a process in which researchers manually select and construct features or covariates that are thought to be relevant for the prediction task. This feature engineering is often based on human intuition, which is inherently subjectively biased and may lead to flawed conclusions about the observation under study. Language models based on neural networks such as Word2Vec (9444008), RNN (liu2016recurrentneuralnetworktext), BERT (devlin2019bertpretrainingdeepbidirectional), by contrast, rely solely on the text itself, bypassing the need for human-driven feature selection. However, these models generally require large amounts of data for training, which poses a significant challenge in social science research, where datasets are often scarce or costly to produce.
Both methods rely heavily on extensive human annotations, making them labor-intensive and challenging for researchers. The time required for hiring, managing, and training labor, coupled with the steep learning curve for model development, renders these methods costly and inefficient to scale. Additionally, these approaches often lack the flexibility to adapt to new data or evolving research questions without task-specific training (terechshenko2020comparison). For instance, a model trained for sentiment analysis to distinguish between positive and negative sentiment cannot be directly applied to a different task, such as topic modeling.
@@
LLMs represent an evolutionary approach to text annotation in political science. A token, like a word in natural language, is the minimum input to an LLM for processing. Auto-regressive LLMs (e.g.touvron2023llamaopenefficientfoundation; geminiteam2024geminifamilyhighlycapable; openai2024gpt4technicalreport), like GPT, operate based on auto-regressive principles, where the model generates text one token at a time, predicting the next token based on the previous ones. Unlike traditional models, LLMs can be treated as black-box tools, accessible via commercial APIs, where users can simply create prompts based on heuristics to direct the model’s output. This capability makes LLMs suitable for a wide range of text annotation and classification tasks without the need for task-specific training. While convenient to use, the classification accuracy of LLMs with simple heuristic prompts is not guaranteed, particularly for less advanced models. The development of LLMs from weak to advanced models has vastly improved their performances. However, the use of advanced LLMs, such as GPT-4, comes with significant practical limitations, including higher cost, slower inference speed, and stricter token usage limit compared to models like GPT-3.5 (Table 1). These challenges pose significant barriers to the widespread application of LLMs in political science. Our proposed framework aims to address these challenges by providing a more accurate, efficient, and cost-effective approach to leveraging LLMs in this field.
@@ 3 Methodology
In the realm of text classification, classic supervised learning approaches necessitate the availability of a large number of annotated examples, upon which models such as SVM and BERT classifiers are trained. Given a task $\mathcal{T}$, a dataset with input-output pairs $\{\textbf{x}_{i},y_{i}\}_{i=1}^{N}$, where $\textbf{x}_{i}$ are the text inputs and $y_{i}$ are the corresponding labels, the optimization objective is to find the model parameters $\theta$ that maximize the likelihood of the correct labels:
$\max_{\theta}\frac{1}{N}\sum_{i=1}^{N}P(y_{i}|\mathbf{x}_{i},\theta)$
These models optimize the likelihood of given labels, adjusting their parameters iteratively to fit the training data. This process, while effective, is often computationally expensive and requires substantial amounts of labeled data to achieve high accuracy.
In contrast, we explore the efficacy of in-context learning (brown2020languagemodelsfewshotlearners), particularly in the context of using auto-regressive large language models (LLM) like Generative Pre-trained Transformer (GPT) for text classification tasks. In-context learning refers to the process where a language model makes predictions based on the context provided by a sequence of input tokens, without updating the model parameters. Instead of updating model parameters $\theta$, the optimization objective of in-context learning for the LLM is to maximize the log-likelihood of the correct class label $y$ given the prompt and query text input $\mathbf{x}$:
$\max_{\text{prompt}}\frac{1}{N}\sum_{i=1}^{N}\log P(y_{i}|\mathbf{x}_{i},\text{prompt};\theta)$
The prompt is natural language text that describes the instructions from humans guiding what the LLM should do. It is usually defined as merely a task description (zero-shot learning) or a task description augmented by a few input-output pairs examples (few-shot learning).
[…]
In zero-shot learning, LLMs make predictions based on common knowledge injected by extensive pre-training conducted by service providers such as OpenAI, Meta, and Google. However, it may not be pertinent enough for a custom user-defined task. Few-shot learning addresses the challenge of injecting special knowledge by leveraging a small number of annotated examples to augment the prompt. The model uses the context provided in the input (e.g., examples of input-output pairs) to make predictions for a new query input. It can be formalized as, given a prompt that includes $k$ ($k<10$) examples of input-output pairs $\{(\mathbf{x}_{1},\mathbf{y}_{1}),(\mathbf{x}_{2},y_{2}),\ldots,(\mathbf{x}_{k},y_{k})\}$ and a new input $\mathbf{x}$, the goal is to predict the corresponding output $y$.
$P(y|\mathbf{x},\text{prompt};\theta)$
Normally, the tradition of designing prompts heavily relies on human heuristics which is usually referred to as prompt engineering. Users create prompts that guide the LLM in producing the desired output in a trial-and-error manner. Although there are practical guidances on how to write prompts, they are far from ideal since users don’t know if the LLM understands the task description or potential pitfalls. In addition, for a task $\mathcal{T}$, only $k$ input-output pairs are pre-selected to augment the prompt for all $N$ examples in the dataset. That is, for every single query text, an LLM sees the same set of examples to make a prediction. However, considering the same set of examples may not be appropriate for all query texts, this simple design cannot maximize the ability of LLM. For example, the sentiment of an example text in Japanese may not be able to guide the sentiment classification of a text in English. Nevertheless, high-performance labeling relies on advanced LLM which are more and more expensive.
To address these challenges and achieve high-performance labeling, we propose an in-context learning framework with three stages: pre-process, in-process, and post-process, as shown in Figure 1. The following sections detail the methodologies and benefits of every stage.
3.1 Pre-process
LLMs face a significant constraint in their context window size, limiting the number of examples that can be used in few-shot learning scenarios. Despite this limitation, there’s often a need to incorporate knowledge from a broader set of examples to enhance the LLM’s performance. To address this challenge, we propose a novel approach: first, we select a pool of $M$ representative examples for human labeling, where $k<M\ll N$. Then, at inference time when calling the commercial API to classify unlabeled query text, we dynamically choose the best $k$ examples to augment the prompt. In the “pre-process” stage, we curate such an exemplar pool for human labeling. This curated pool serves two crucial purposes: it allows the LLM to infer the underlying rules used for labeling, and it enables adaptive selection of the most relevant examples for each query text. By creating this initial pool of human-annotated examples, we lay the foundation for more effective and efficient use of the LLM in subsequent stages, balancing the need for comprehensive knowledge injection with the constraints of the model’s context window. In this stage, we prepare this exemplar pool, and here are the steps.

### v2

# Enhancing LLM-Based Text Classification in Political Science: Automatic Prompt Optimization and Dynamic Exemplar Selection for Few-Shot Learning

@@ Abstract
Large language models (LLMs) offer substantial promise for text classification in political science, yet their effectiveness often depends on high-quality prompts and exemplars. To address this, we introduce a three-stage framework that enhances LLM performance through automatic prompt optimization, dynamic exemplar selection, and a consensus mechanism. Our approach automates prompt refinement using task-specific exemplars, eliminating speculative trial-and-error adjustments and producing structured prompts aligned with human-defined criteria. In the second stage, we dynamically select the most relevant exemplars, ensuring contextually appropriate guidance for each query. Finally, our consensus mechanism mimics the role of multiple human coders for a single task, combining outputs from LLMs to achieve high reliability and consistency at a reduced cost. Evaluated across tasks including sentiment analysis, stance detection, and campaign ad tone classification, our method enhances classification accuracy without requiring task-specific model retraining or extensive manual adjustments to prompts. This framework not only boosts accuracy, interpretability and transparency but also provides a cost-effective, scalable solution tailored to political science applications. An open-source Python package (PoliPrompt) is available on GitHub.
@@
text classification, sentiment analysis, stance detection, in-context learning, automatic prompt optimization, large language models (LLMs)
@@ 1 Introduction
Text data is a key resource in political science, where text classification tasks—such as stance detection, sentiment analysis, and topic classification—are widely used to study political discourse and public opinion. For example, sentiment analysis of campaign ads can track trends in messaging tone, helping researchers understand their impact on voter perception (fowler2021political). In authoritarian contexts, text classification sheds light on public opinion about government actions, as demonstrated by pan2018concealing, who categorize online posts in China by support or criticism of local governance, and by king2013censorship, who analyze censorship patterns in Chinese social media. Text classification also plays an essential role in legislative studies. grimmer2013text) show how categorizing legislative texts and party manifestos helps quantify ideological shifts over time, while slapin2008scaling develop a model to track party positions by analyzing speeches in parliamentary debates. On social media, classification has been used to examine political polarization and public engagement. bond2015quantifying analyze social media topics to assess voter engagement, and barbera2015understanding uses Twitter data to reveal ideological clustering. Collectively, these studies demonstrate how text classification is a common method in political science for analyzing complex data and understanding political behavior.
Political scientists have long relied on training-based methods for text classification, including Statistical Language Models (SLMs) and, more recently, Pre-trained Language Models (PLMs) (8950616; 9079815; 1174243; devlin2019bertpretrainingdeepbidirectional; sun2020finetuneberttextclassification). SLMs, often paired with classifiers like Support Vector Machines (SVMs), capture basic word patterns but require extensive manual feature engineering, making them labor-intensive. Importantly, SLMs are highly task-specific: a model trained for one purpose, such as topic classification, must be retrained for a different task like sentiment analysis or for application on new types of text, such as social media posts versus news articles. PLMs, like BERT, offer advancements by generating context-aware word representations that require less manual feature engineering and capture more subtle relationships between words.
In recent years, Large Language Models (LLMs) have revolutionized text classification by enabling flexible and efficient learning without extensive task-specific training. A key mechanism driving this capability is in-context learning (ICL), where LLMs perform classification using only a well-crafted prompt, eliminating the need for extensive task-specific training or fine-tuning on labeled data. [^1] A prompt typically consists of four main components: (1) task description, which provides context for the model, (2) few-shot examples, which guide the model’s responses based on prior cases, (3) the query, where the model generates predictions based on the given input question or context, and (4) the prefix, which serves as an additional guiding phrase or instruction to shape the model’s output format (brown2020languagemodelsfewshotlearners). 222 A detailed comparison between in-context learning and fine-tuning is presented in Table A2 in the appendix, highlighting their respective advantages, limitations, and use cases.This allows LLMs to be applied seamlessly across diverse classification tasks (kaplan2020scaling). While ICL offers significant advantages, its effectiveness heavily depends on handcrafted prompts which introduce several challenges. Manually designed prompts through prompt engineering remain an ad-hoc and trial-and-error process, requiring domain expertise to craft effective inputs. Despite its accessibility, prompt engineering exhibits black-box characteristics, as the exact way in which LLMs interpret and respond to prompts is often opaque. Small modifications in wording or structure can lead to unpredictable variations in output, making it difficult to establish clear causal relationships. Additionally, the effectiveness of a prompt is often model-specific, meaning that a well-performing prompt for one LLM may not generalize across different architectures. These limitations underscore the need for systematic and automated prompt optimization to improve consistency and performance.
Our proposed three-stage LLM-based framework addresses key limitations of existing text classification approaches by introducing automatic prompt optimization, dynamic exemplar selection, and a consensus mechanism. These methods, which have been adopted and adapted from computer science, are systematically integrated into political science research to enhance LLM performance. Our contributions include introducing these concepts to political science, integrating them into a unified framework, providing a simple and accessible implementation, and streamlining techniques by excluding unnecessary complexity for this task. This framework directly addresses critical challenges in LLM-based text classification. Automatic prompt optimization overcomes the trial-and-error nature of prompt engineering by systematically refining prompts, reducing reliance on manual tuning, and improving consistency. Dynamic exemplar selection mitigates the issue of LLM sensitivity to few-shot examples [^3] Figure A1 presents a comparison of zero-shot and few-shot prompting, demonstrating how structured inputs guide LLM decision-making. by ensuring that the most relevant examples are selected for each classification task, leading to more stable and reliable outputs. The consensus mechanism improves classification reliability by reducing variability in LLM-generated responses, addressing inconsistencies, and improving agreement across model outputs. [^4] A summary of how our framework differs from current practices in political science is provided in Table A1 in the appendix.
Existing prompt engineering relies on intuition-driven trial and error, offering little feedback on where and how prompts fail, making generalization difficult. To address this, we propose using automatic prompt optimization in the first stage, which systematically refines prompts using human-labeled exemplars rather than subjective adjustments. Our data-driven approach curates a small set of labeled examples (typically ¡100) to generate an optimized prompt tailored to task requirements. This method minimizes guesswork, enhances replicability, and improves classification accuracy over prompt engineering. The effectiveness of our method is demonstrated through quantitative and qualitative evaluations.
The second stage, dynamic exemplar selection, tailors the prompt exemplars to the most contextually relevant cases for each query rather than relying on a fixed set. This adaptation enhances classification accuracy by making the model’s inference process more context-sensitive. Importantly, the selected exemplars are not only the most relevant to the query text but also representative of the dataset as a whole, ensuring that the model learns from diverse and structurally significant examples. This is particularly crucial when dealing with complex data structures, where certain underrepresented patterns or relationships could otherwise be overlooked. By maintaining both query relevance and dataset representativeness, this method strengthens the model’s ability to handle subtle conceptual boundaries more effectively and generalize across diverse classification tasks.
In the final stage, our framework employs a consensus mechanism in which two LLMs independently generate predictions, and a third LLM serves as a judge to resolve any disagreements. Rather than simply aggregating outputs, the judge model evaluates the reasoning behind each prediction—often through a Chain-of-Thought (CoT) process—and determines the more accurate classification. This approach mirrors human annotation workflows, where a third reviewer adjudicates conflicts between two coders, thereby increasing reliability and interpretability. By incorporating a structured resolution step, the LLM-as-judge mechanism reduces classification noise and improves robustness, offering a scalable solution for tasks that require high precision and consistency.
We apply our framework to three classification tasks: sentiment analysis, stance detection, and multi-category campaign ad tone classification. Our results show that automatic prompt optimization provides a clear alternative to trial-and-error prompt engineering, offering structured guidance to distinguish commonly conflated concepts—such as sentiment versus stance—without requiring new model training. In the tone classification task, our consensus mechanism further enhances reliability by mimicking multi-coder setups and flagging labeling inconsistencies, a key feature for ensuring data quality in subjective tasks. These applications demonstrate the versatility and robustness of our method for a wide range of political science classification problems.
@@ 2 Literature Review
The evolution of BERT’s ”pre-training and fine-tuning” paradigm led to the creation of Large Language Models (LLMs), a subset of Pre-trained Language Models (PLMs) characterized by an immense scale of parameters and training data. Following a principle known as the scaling law (kaplan2020scaling), researchers observed that increasing both the size of the model (in terms of parameters) and the amount of training data leads to significant improvements in the model’s overall performance across a wide variety of tasks. A key outcome of scaling is what are termed emergent abilities—complex skills such as interpreting complex instructions and performing multi-step reasoning (wei2022emergent). Among these abilities is in-context learning, which allows LLMs to interpret and execute tasks based solely on prompt instructions, eliminating the need for task-specific re-training or fine-tuning (dong2022survey). This capability enables LLMs like GPT-3 to leverage their pre-trained knowledge directly for new tasks, distinguishing them from earlier models that rely heavily on additional training to adapt to different applications.
One subcategory of in-context learning is zero-shot learning (brown2020languagemodelsfewshotlearners; radford2019language), where LLMs can perform tasks based purely on a description of the task, without requiring any labeled exemplars. For instance, GPT-3 can classify the sentiment of a sentence (positive or negative) based solely on a prompt describing the task. Additionally, LLMs support few-shot learning (brown2020languagemodelsfewshotlearners; min-etal-2022-rethinking), where the inclusion of a handful of exemplars in the prompt enables the model to better understand and generalize the task, further improving accuracy. This flexibility makes LLMs especially valuable in fields like political science, where labeled data is often scarce. By generating text token by token, they maintain coherent context across long passages, enabling them to handle a range of tasks based on a single prompt. This capability provides political scientists with a powerful, adaptable tool for text analysis without the need for fine-tuning, positioning LLMs as a resource-efficient alternative to traditional methods (radford2019language).
LLMs offer powerful capabilities for text classification, yet their effectiveness often depends heavily on well-designed prompts, a reliance that has led to the labor-intensive practice of prompt engineering (khattab2023dspy). Research applying LLMs to social science tasks has underscored this challenge, showing that LLMs alone may not consistently achieve sufficient accuracy without tailored prompt adjustments or additional human oversight. For example, ChatGPT, when used to classify social media sentiment on HPV vaccination, exhibited higher accuracy for anti-vaccination messages but struggled with pro-vaccination sentiment, particularly for longer formats, indicating the model’s sensitivity to both content type and length (kim2024accuracy). Similarly, efforts to classify policy documents into specific issue categories using GPT-3.5 and GPT-4 revealed that LLMs achieved only moderate accuracy without human intervention—ranging from 65–83%—and required significant manual oversight for contentious or complex cases (gunes2023multiclass). These studies suggest that while LLMs offer promising automation potential, optimal results often require extensive prompt customization and careful selection of exemplar data.
A key characteristic of in-context learning (ICL) is its ability to optimize the context provided before generation rather than modifying model parameters. Instead of requiring additional training, ICL conditions the model through carefully structured prompts that influence its responses. Studies have shown that well-structured few-shot prompts can significantly improve LLM performance in text classification (min-etal-2022-rethinking), reinforcing the importance of effective context design in achieving reliable results.
The prompt engineering workflow (illustrated in Figure A2) involves three main stages: creating an initial prompt, testing the prompt on the model to generate results, and then verifying those results extensively. If the output yields low accuracy, the user re-enters a Revise and Try loop, repeatedly adjusting the prompt without clear feedback on how each change impacts model behavior. This blind, trial-and-error process can be time-consuming and uncertain, as users often lack systematic guidance for prompt refinement. Once a satisfactory level of accuracy is achieved, the final prompt is established for consistent application.
Despite its utility, prompt engineering has notable limitations. Ensuring that a prompt is both clear and accurately interpreted by the model remains challenging, as users often rely on summary metrics like F1 scores to assess quality. However, these metrics offer little guidance on how to improve prompts, leading to an iterative cycle of speculative revisions without assurance of improved performance. Moreover, prompt effectiveness tends to be model-specific; a prompt optimized for one LLM may perform poorly on another, limiting generalizability across different models and data contexts (khattab2023dspy). Although in-context learning increases adaptability by using exemplars within the prompt, it often relies on static exemplars. As the model encounters varied or new text data, these static exemplars may not align well, potentially reducing output accuracy.
Recent work has introduced automatic prompt optimization methods like DSPy (khattab2023dspy), APE (zhou2022large), OPRO (yang2024largelanguagemodelsoptimizers), and EvoPrompt (guo2024optimizersbetteronellm), which aim to improve prompt performance by searching for high-performing strings. For instance, DSPy organizes tasks into step-by-step pipelines using modules called ”teleprompters” that fine-tune model behavior at each step. These multi-stage approaches work well for tasks requiring structured reasoning or complex retrieval, like multi-hop question answering, but are often computationally intensive and may not translate effectively to political science text classification. DSPy’s complex decomposition and reliance on advanced models make it resource-demanding, limiting its practicality for more context-sensitive text analysis, like assessing sentiment or stance in political texts, which often requires careful contextual interpretation and is common in smaller-scale research projects.
To meet the specific demands of political science text analysis, we introduce a novel approach incorporating automatic prompt optimization, dynamic exemplar selection, and a consensus mechanism, which we will explain in detail in the following methodology section.
@@ 3 Methodology
3.1 Data Preprocessing
Before applying automatic prompt optimization, we first process and structure the dataset to enable efficient exemplar selection. This preparation stage includes two key steps: converting texts into embeddings and reducing feature dimensionality. These transformations help organize the dataset to support dynamic exemplar selection while optimizing use of the LLM’s context window.